OpenStudy (anonymous):

a wise man once said, 400 reduced by 4 times my age is 188

OpenStudy (jdoe0001):

what was his name? don't leave us in suspense!

OpenStudy (jdoe0001):

400 - 4 times (his age) = 188 say his age is "x" 400 - 4x = 188 solve for "x"

OpenStudy (anonymous):

Ok his name is jdoe0001

OpenStudy (jdoe0001):

ohhh that guy.... well, I think I know what that is hehe

OpenStudy (accidentalaichan):

400-4x=188 -440 -400 -4x = -12

OpenStudy (accidentalaichan):

*-212

OpenStudy (anonymous):

Ok thx

OpenStudy (accidentalaichan):

Now, divide -212 by -4 and there's your answer.